OpenStudy (anonymous):

An open top rectangular box must have a volume of 4500 cu ft. If the bottom is a rectangle with length is twice the width, what dimensions would minimize total surface area of the box?

OpenStudy (anonymous):

V = l w h SA = 2(lw + lh + wh) l = 2w V = 2 w w h SA = 2(2w w + 2wh + wh) 4500 = 2 w w h h = 4500 / 2 w w SA = 2(2 w w + 2 w (4500 / 2 w w ) + w (4500 / 2 w w )) SA = 4*w w+13500*w w w SA' = ... Minimize it simply